OpenStudy (anonymous):

Anyone know much about solving linear equations with the method developed in India? x(sqrd)+3x-10=0 x(sqrd)+3x=10 4x(sqrd)+12x=40 4x(sqrd)+12x+9=40+9 4x(sqrd)+12x+9=49 2x+3=-+7 I dont understand how the example takes the square root of both sides on the last part of the example. Where does the 12x go?

OpenStudy (anonymous):

Quadratic Equations....sorry

OpenStudy (anonymous):

Use the formula \[(a+b)^2 = a^2 +2ab +b^2\] \[2*a*b = 2*(2x)*(3)\]